40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 5 11:33:10 2014, Koji, Summary, LSC, 2 arm ALS->LSC transition - road map 
    Reply  Wed Mar 5 19:15:39 2014, Jenne, Summary, LSC, ALS offset moving script modified 
    Reply  Wed Mar 5 22:32:21 2014, manasa, Update, LSC, Stuck at step 2 
       Reply  Thu Mar 6 11:15:32 2014, Koji, Summary, LSC, Stuck at step 2 
    Reply  Fri Mar 7 00:43:34 2014, manasa, Update, LSC, ALS C&D locked (on MC2 and ETMs) and IR resonance obtained ALS_MC2CARM.pdf
Message ID: 9702     Entry time: Fri Mar 7 00:43:34 2014     In reply to: 9691
Author: manasa 
Type: Update 
Category: LSC 
Subject: ALS C&D locked (on MC2 and ETMs) and IR resonance obtained 

[EricQ, Manasa]

ALS common locked by actuating on MC2 and ALS Differential locked by actuating on ETMX and ETMY (Stable lock acquired for over an hour).

Common and Differential offsets were swept to obtain IR resonance in both the arms (arms stayed on resonance for over 15 minutes).

Procedure:

1. Configured LSC settings to allow locking using ALS error signals.

2. Locked common and differential using ALS error signals

Aux matrix
              ALSX    ALSY
------------------------------
XARM    1            -1
YARM    1              1
-----------------------------
X arm servo settings:
FIlters: FM1, FM5, FM6, FM7, FM9
Gain = -8.0

Y arm servo settings:
Filters: FM1, FM5, FM6, FM7, FM9
Gain = +8.0

Out matrix
    XARM    YARM
------------------------
ETMX    1    0
ETMY    0    1
------------------------

3. Transitioned CARM control output to actuate on MC2 instead of ETMX

SUS-MC2_LSC servo gain = 1.0
The transition was done in very small steps : actuating on MC2 in -0.01 steps at the outmatrix upto -1.0 while reducing the ETMX actuation to 0 simultaneously.

DARM still stayed locked only with actuation on ETMY.

4. Transitioned DARM control to ETMX and ETMY.

Used ezcastep to step up DARM control (Y arm output) actuation on ETMX and step down the actuation on ETMY.

Final output matrix
    Xarm    Yarm
-----------------------
ETMX      0    -0.5
ETMY      0     0.5
MC2    -1.0      0
-----------------------

Noise plot in attachment.

5. Finding arm resonance

Used ezcastep to gradually build up offsets in CARM (LSC-XARM_OFS) to find IR resoance in one arm (Y arm).
Introducing a small (order of 0.5) DARM offset (LSC-YARM_OFS) shifted the Y arm off-resonance.
Used CARM offset to get back the Y arm to resonance.
Changing CARM and DARM offsets alternately while tracking the Y arm resonance got us to a point where we had both the arms resonating for IR.

6. At this point the MC decided to give up and we lost lock.

Notes:
1. We found that the WFS2 YAW output filterbank had the output switched OFF (probably accidentally by one of us). This was reenabled. Please be careful while manually turning ON and OFF the MC WFS servos.

Attachment 1: ALS_MC2CARM.pdf  82 kB  Uploaded Fri Mar 7 13:31:45 2014  | Hide | Hide all
ALS_MC2CARM.pdf
ELOG V3.1.3-